During a promotional event,a sporting goods store gave a free T-shirt to every 8th customer and a free water bottle to every 10th customer.Which customer was the first to get a free T-shirt and a free water bottle?
1
Expert's answer
2019-09-06T09:21:00-0400

The number of customers is divisible by 8 and divisible by 10. Since 8=2*2*2, 10=2*5, we have 2*2*2*5=40 is the least number with such a property. Thus, the 40th customer was the first to get a free T-shirt and a free water bottle.